# Enhancing the Accuracy of Spatio-Temporal Models for Wind Speed Prediction by Incorporating Bias-Corrected Crowdsourced Data

### v1

@@ Abstract
Accurate high-resolution spatial and temporal wind speed data is critical for estimating the wind energy potential of a location. For real-time wind speed prediction, statistical models typically depend on high-quality (near) real-time data from official meteorological stations to improve forecasting accuracy. Personal weather stations (PWS) offer an additional source of real-time data and broader spatial coverage than official stations. However, they are not subject to rigorous quality control and may exhibit bias or measurement errors. This paper presents a framework for incorporating PWS data into statistical models for validated official meteorological station data via a two-stage approach. First, bias correction is performed on PWS wind speed data using reanalysis data. Second, we implement a Bayesian hierarchical spatio-temporal model that accounts for varying measurement error in the PWS data. This enables wind speed prediction across a target area, and is particularly beneficial for improving predictions in regions sparse in official monitoring stations. Our results show that including bias-corrected PWS data improves prediction accuracy compared to using meteorological station data alone, with a 7% reduction in prediction error on average across all sites. The results are comparable with popular reanalysis products, but unlike these numerical weather models our approach is available in real-time and offers improved uncertainty quantification.
@@ 1 Introduction
In this paper we investigate the benefits of incorporating crowdsourced wind data — a previously untapped source of information — into spatio-temporal models for predicting wind speeds at unobserved locations in Ireland. While the focus in this paper is the use of wind speed data for the evaluation of the potential renewable energy resource at a specific location, as well as management of installed infrastructure, accurate data on wind speeds at precise temporal and spatial scales is crucial in a number of application areas. The include: tracking storm conditions; the spread of wild fires (Beer,, 1991), and modelling air pollution (Sahoo et al.,, 2023). To combat the effects of climate change, improve air quality and achieve energy independence, EU countries aim to accelerate the transition to renewable electricity generation (European Union,, 2022). The EU has set a target of 45% of electricity production to be generated from renewable sources by 2030 (European Union,, 2022). Similarly, the Irish government’s Climate Action Plan 2024 (CAP24, Department of the Environment, Climate and Communications, (2024)) has set a target of 80% of national electricity production to be generated from renewable sources by 2030. Wind power is Ireland’s primary renewable resource, contributing 35% of Irish electricity in 2023 (Wind Energy Ireland,, 2023). In CAP24 the Irish government has set a target of 9GW of onshore wind energy by 2030, and at least 5GW of off-shore wind energy  (Department of the Environment, Climate and Communications,, 2024). Currently, the Republic of Ireland has approximately 4.8GW of onshore wind energy installed, highlighting a 4.2GW gap (Wind Energy Ireland,, 2025). The peak daily electricity demand ranges between 4.5 and 5.5 GW, with highs of more than 6 GW during cold weather.
@@
The novelty of our approach is to explore the the incorporation of previously unstudied crowdsourced weather observations to increase the quantity and spatial resolution of real-time wind speed data in the Republic of Ireland. These models enable the estimation of wind speeds at unobserved locations, an important variable for wind resource assessment. We focus on data sources available in near real-time, with the goal of improving operational capabilities and laying the groundwork for future research in nowcasting techniques. Crowdsourcing for data collection has been widely used in other fields, such as astronomy (Raddick et al.,, 2009), ecology (Mengersen et al.,, 2017), and air pollution monitoring (Bonas and Castruccio,, 2023). In the context of weather, crowdsourced data refers to individuals uploading weather measurements using personal weather stations (PWS), which has gained prominence with the advent of lower-cost sensors and internet connectivity. One such project is the weather observation website (WOW, Weather Observation Website, 2011a ) in collaboration with Ireland’s official meteorological organisation, Met Éireann and other national meteorological services. There is extensive information available for the Republic of Ireland, the United Kingdom and the Netherlands at wow.met.ie. In these countries, there are more crowdsourced stations than official meteorological ones. For example, approximately 100 personal stations have uploaded data to the WOW website in the Republic of Ireland, and over 10,000 are available globally. This rich dataset has the potential to enhance high-resolution wind resource estimation and improve short-term forecasting. However, the lack of quality control for these data sources, along with possible biases and noisy observations often impedes their practical use. This is a common issue in crowdsourced data across all fields (Aceves-Bueno et al.,, 2017). For weather there has been a significant body of research on utilising crowdsourced data with potential quality issues  (Chakraborty et al.,, 2020; Chen et al.,, 2018), but the majority has been on variables such as temperature or air pressure, which demonstrate less local variability.
@@
Wind data presents unique challenges due to its high variability and dependence on localized factors such as obstacles and the altitude of anemometers (wind speed measurement instruments). Some studies have begun addressing these challenges. For example, Droste et al., (2020) analysed crowdsourced wind data in the Netherlands using descriptive statistics to examine the distribution of wind speeds at meteorological and PWS, in particular highlighting issues with devices underestimating true wind speeds. They applied a linear correction method based on co-located meteorological and PWS (Droste et al.,, 2020). While this improves the accuracy of the crowdsourced stations, in our dataset, meteorological and crowdsourced stations are not co-located.  Chen et al., (2021) filtered crowdsourced stations by applying initial quality control measures such as examining correlation of the crowdsourced station with the nearest meteorological station, and ensuring temporal consistency with the measurements. With the remaining stations, bias correction was performed using a quantile mapping approach, with quantiles interpolated from meteorological stations. However, Ireland contains both low number of meteorological weather stations, and a climate with highly variable wind speeds (Met Éireann, 2025b, ).
@@
This paper is structured as follows: in Section 2 an overview of the datasets used is provided and quality control measures are discussed. The numerical weather models used for bias correction are also introduced. In Section 3, a bias correction is performed for the PWS data sources using quantile transformations and incorporating reanalysis data. In Section 4 spatial and spatio-temporal models are introduced along with extensions to mitigate noisy data. In Section 5 the models are applied to Irish wind data, and the predictive improvement of incorporating crowdsourced data are evidenced.
@@ 2 Data
The wind data sources in this article consist of official meteorological data, citizen science or crowdsourced wind data in the form of personal weather stations, and reanalysis data from numerical weather simulation models. We exclusively focus on data from the Republic of Ireland, an island located in the east Atlantic. Its climate is strongly influenced by the ocean, with the prevailing wind between the south and west, and wind strongest in coastal locations (Met Éireann, 2025b, ). An overview of each data source is provided in the following.
@@ 2.1 Met Éireann Data
As Ireland’s national meteorological body, Met Éireann maintains a network of quality-controlled weather stations, adhering to the standards set by the World Meteorological Organization (Oke,, 2018). Meteorological station data sourced from Met Éireann is regarded as ground truth in this study due to the high-quality measurement devices used and the optimal siting of instruments, ensuring no obstacles in the vicinity of measurement stations (Met Éireann, 2025a, ). Wind speed measurements are taken at a standard height of 10m in open locations, ensuring no obstructions within a 100m radius (Met Éireann, 2025b, ). The data is available in one minute intervals for 18 stations, and at hourly level for a further 5 stations. For consistency across all stations, including the PWS data introduced in Section 2.2, the data is aggregated to an hourly level by averaging sub-hourly measurements over the preceding hour. Aggregating high-frequency wind measurements is a standard practice in wind resource studies, as it balances temporal resolution with noise reduction and compatibility with other datasets (Torres et al.,, 2005). Figure 1 shows the location of the stations, and the median and inter-quartile range of wind speeds over a full year. Although there are only 23 weather stations which measure wind speeds in Ireland—a relatively small number—they are distributed across the country, ensuring good geographic coverage. In general, median wind speeds and wind speed variability are highest in coastal locations, particularly the west coast.
Wind speeds also exhibit a daily or diurnal pattern, as demonstrated in Figure 2, which presents the average daily wind speeds for each hour of the day, at four station locations as an example case study. This shows higher wind speeds during the afternoon with lower wind speeds at night. This pattern is commonly seen in wind speeds and attributed to the effects of solar heating (Dai and Deser,, 1999).
@@
Figure 2: Average wind speeds by hour at four meteorological stations across a full year (2024). The four sample stations are numbered on the map of median wind speeds in Figure 1.
@@ 2.2 Personal Weather Stations
The personal weather station (PWS) data is obtained from Weather Observation Website, 2011a , which provides measurements from stations set up by individual users. We selected this dataset because it is hosted on a single platform alongside meteorological station data and includes supporting information on station setup. Other sources of crowdsourced weather data include commercial platforms such as Weather Underground, (2025) and Netatmo, (2025), which could be considered in future research. The data from each station is sent digitally to the website which displays weather data and metadata on all stations. Unlike meteorological stations, there is no consistency of equipment or setup, nor is there any quality control. The website provides information on the location in addition to a broad classification of the site attributes. These fall under three broad classifications, which are based on grading schemes from meteorological bodies (Weather Observation Website, 2011b, ), along with other stations which have not been classified. The three main classifications are described in Table 1:
@@
A Wind sensors calibrated within last 10 years, mounted 10m above the ground on mast or pole, with no obstructions within 100m.
@@
Temporal resolution of the stations can vary, although they are generally at most hourly, and frequently contain data on a sub hourly resolution, many of which are on a 5 minute resolution.
Figure 3 shows the distribution of wind speeds across different measurement sources, and highlights that wind speeds at PWS are generally lower when compared to meteorological stations, a similar finding to previous studies on crowdsourced data (e.g. Chen et al.,, 2021). This may be due to the high dependency of wind measurements on the placement of equipment. Unlike weather variables such as temperature or air-pressure, local obstacles such as buildings, trees or hedges can have a significant effect on local wind speeds. Meteorological stations typically place anemometers at 10m heights, whereas the heights of the PWS measurements are assumed to be at less than 10m unless the station is in class A.
Figure 3: Violin plot showing the wind speed distribution for each group of stations in the Republic of Ireland. A,B and C are different classes of PWS. Met are meteorological weather stations, and U is PWS with unknown quality. The black diamonds denote the median of each group.
Consequently, directly including PWS wind speeds in a model, be it statistical or a physical weather model, would most likely lead to significant inaccuracies, in particular underestimation of wind speeds (Chen et al., (2021)). To highlight the potential usefulness of crowdsourced stations we select examples where we have a meteorological station and PWS in close proximity($<$15km) to each other, presented in Figure 4.
Figure 4: Location of meteorological stations in blue, and PWS in red. Four groups have been highlighted and sample time series for each are shown.
@@
(a) 1. Mace Head (Met) and Glynsk (PWS). Pearson correlation = 0.86
(b) 2. Sherkin Island (Met) and Skibbereen (PWS). Pearson correlation = 0.76
(c) 3. Dublin Airport (Met) and Swords (PWS). Pearson correlation = 0.79
(d) 4. Oak Park (Met), Ballyhide (PWS). Pearson correlation = 0.83
@@
Figure 5: A comparison of nearby meteorological (Met) stations (in blue) and PWS (in red) at four locations around Ireland over a randomly selected three day period. Each caption contains the Pearson correlation coefficient between the two time series.
@@
Although Figure 5 illustrates that PWS tend to underestimate the true wind speeds, there appears to be a correlation between the time series of the neighbouring stations. To compare measurements from nearby stations on a common scale, we compute the empirical cumulative distribution function (ECDF) for each station, which expresses each observation as its percentile rank within that station’s distribution. This transform all values to a [0,1] scale. Figure 6 shows good agreement between neighbouring stations, which exhibit similar trends over time and comparable percentile ranks at each time point.
(a) c Mace Head (Met) and Glynsk (PWS). Pearson correlation = 0.85
(b) 2. Sherkin Island (Met) and Skibbereen (PWS). Pearson correlation = 0.78
(c) 3. Dublin Airport (Met) and Swords (PWS). Pearson correlation = 0.80
(d) 4. Oak Park (Met) and Ballyhide (PWS). Pearson correlation = 0.83
Figure 6: A comparison of probabilities from the empirical cumulative density functions (ECDF) of nearby meteorological stations and PWS at four locations around Ireland. The ECDFs shown here are computed from the entire dataset, while the plots display a three-day sample period.
@@
In general, PWS data appears to be correlated with nearby meteorological stations. However, as PWS can vary widely in reliability and data quality, we apply a series of quality assurance checks, and identify and remove stations that do not meet these checks. Bias correction methods for addressing remaining systematic differences are described in Section 3.
@@
First, we exclude PWS stations with excessive missing data. Only stations with at least 90% non-missing wind speed observations are retained, a criteria used in Lenzi and Genton, (2020). In addition, stations that do not upload data in near real-time are excluded, since applications in renewable energy require timely updates.
@@
Second, we assess data quality using empirical correlations. Specifically, we compute Spearman’s rank correlation coefficients between the wind time series at each station (Figure 7). Spearman’s method is robust to non-linear relationships and captures consistency in wind speed rankings rather than raw values—important when stations may exhibit non-linear biases. As expected, meteorological stations display the highest pairwise correlations. However, the trend of decreasing correlation with distance is also seen for Met-PWS and PWS–PWS station pairs.
@@
To remove unreliable stations, we extend the approach of Chen et al., (2021), who discard stations with correlation below 0.5 relative to the nearest meteorological station. In our case, a PWS is retained only if at least five of its nearest neighbours (which may be meteorological or personal) exhibit a Spearman correlation greater than 0.5. This threshold is chosen as all meteorological stations meet this criteria, indicating it is a reasonable benchmark.

### v2

@@ Abstract
Accurate high-resolution spatial and temporal wind speed data is critical for estimating the wind energy potential of a location. For real-time wind speed prediction, statistical models typically depend on high-quality (near) real-time data from official meteorological stations to improve forecasting accuracy. Personal weather stations (PWS) offer an additional source of real-time data and broader spatial coverage than official stations. However, they are not subject to rigorous quality control and may exhibit bias or measurement errors. This paper presents a framework for incorporating PWS data into statistical models for validated official meteorological station data via a two-stage approach. First, bias correction is performed on PWS wind speed data using reanalysis data. Second, we implement a Bayesian hierarchical spatio-temporal model that accounts for varying measurement error in the PWS data. This enables wind speed prediction across a target area, and is particularly beneficial for improving predictions in regions sparse in official monitoring stations. Our results show that including bias-corrected PWS data improves prediction accuracy compared to using meteorological station data alone, with a 5% reduction in prediction error on average across all sites. The results are comparable with popular reanalysis products, but unlike these numerical weather models our approach is available in real-time and offers improved uncertainty quantification.
This is the Author Accepted Manuscript of the article: Organ, E., Upton, M., Allard, D., Benoit, L., and Sweeney, J. (2025), “Enhancing the Accuracy of Spatio-Temporal Models for Wind Speed Prediction by Incorporating Bias-Corrected Crowdsourced Data,” Environmetrics. The final Version of Record is available at:
https://doi.org/10.1002/env.70069.
This manuscript is made available under the Creative Commons Attribution 4.0 International (CC BY 4.0) licence: https://creativecommons.org/licenses/by/4.0/.
@@ 1 Introduction
In this paper we investigate the benefits of incorporating crowdsourced wind data — a previously untapped source of information — into spatio-temporal models for predicting wind speeds at unobserved locations in Ireland. While the focus in this paper is the use of wind speed data for the evaluation of the potential renewable energy resource at a specific location, as well as management of installed infrastructure, accurate data on wind speeds at precise temporal and spatial scales is crucial in a number of application areas. The include: tracking storm conditions (Lagerquist et al.,, 2017), the spread of wild fires (Beer,, 1991), and modelling air pollution (Sahoo et al.,, 2023). To combat the effects of climate change, improve air quality and achieve energy independence, EU countries aim to accelerate the transition to renewable electricity generation (European Union,, 2022). The EU has set a target of 45% of electricity production to be generated from renewable sources by 2030 (European Union,, 2022). Similarly, the Irish government’s Climate Action Plan 2024 (CAP24, Department of the Environment, Climate and Communications, (2024)) has set a target of 80% of national electricity production to be generated from renewable sources by 2030. Wind power is Ireland’s primary renewable resource, contributing 35% of Irish electricity in 2023 (Wind Energy Ireland,, 2023). In CAP24 the Irish government has set a target of 9GW of onshore wind energy by 2030, and at least 5GW of off-shore wind energy (Department of the Environment, Climate and Communications,, 2024). Currently, the Republic of Ireland has approximately 4.8GW of onshore wind energy installed, highlighting a 4.2GW gap (Wind Energy Ireland,, 2025). The peak daily electricity demand ranges between 4.5 and 5.5 GW, with highs of more than 6 GW during cold weather.
@@
The novelty of our approach is to explore the incorporation of previously unstudied crowdsourced weather observations to increase the quantity and spatial resolution of real-time wind speed data in the Republic of Ireland. We focus on data sources available in near real-time, with the goal of improving operational capabilities and laying the groundwork for future research in nowcasting techniques. Crowdsourcing for data collection has been widely used in other fields, such as astronomy (Raddick et al.,, 2009), ecology (Mengersen et al.,, 2017), and air pollution monitoring (Bonas and Castruccio,, 2023). In the context of weather, crowdsourced data refers to individuals uploading weather measurements using personal weather stations (PWS), which has gained prominence with the advent of lower-cost sensors and internet connectivity. One such project is the weather observation website (WOW, Weather Observation Website, 2011a ) in collaboration with Ireland’s official meteorological organisation, Met Éireann and other national meteorological services. There is extensive information available for the Republic of Ireland, the United Kingdom and the Netherlands at wow.met.ie. In these countries, there are more crowdsourced stations than official meteorological ones. For example, approximately 100 personal stations have uploaded data to the WOW website in the Republic of Ireland, and over 10,000 are available globally. This rich dataset has the potential to enhance high-resolution wind resource estimation and improve short-term forecasting. However, the lack of quality control for these data sources, along with possible biases and noisy observations often impedes their practical use. This is a common issue in crowdsourced data across all fields (Aceves-Bueno et al.,, 2017). For weather there has been a significant body of research on utilising crowdsourced data with potential quality issues (Chakraborty et al.,, 2020; Chen et al.,, 2018), but the majority has been on variables such as temperature or air pressure, which demonstrate significantly less local variability than wind.
@@
Wind data presents unique challenges due to its high variability and dependence on localized factors such as obstacles and the altitude of anemometers (wind speed measurement instruments). Some studies have begun addressing these challenges. For example, Droste et al., (2020) analysed crowdsourced wind data in the Netherlands using descriptive statistics to examine the distribution of wind speeds at meteorological and PWS, in particular highlighting issues with devices underestimating true wind speeds. They applied a linear correction method based on co-located meteorological and PWS (Droste et al.,, 2020). While this improves the accuracy of the crowdsourced stations, in our dataset, meteorological and crowdsourced stations are not co-located. Chen et al., 2021c filtered crowdsourced stations by applying initial quality control measures such as examining correlation of the crowdsourced station with the nearest meteorological station, and ensuring temporal consistency with the measurements. With the remaining stations, bias correction was performed using a quantile mapping approach, with quantiles interpolated from meteorological stations. However, Ireland contains both low number of meteorological weather stations and a climate with highly variable wind speeds (Met Éireann, 2025b, ), which makes quantile interpolation uncertain.
@@
This paper is structured as follows: in Section 2 an overview of the datasets used is provided and quality control measures are discussed. The numerical weather models used for bias correction are also introduced. In Section 3, a bias correction is performed for the PWS data using quantile transformations and incorporating reanalysis data. In Section 4 spatial and spatio-temporal models are introduced along with extensions to mitigate noisy data. In Section 5 the models are applied to Irish wind data, and the predictive improvements of incorporating crowdsourced data are evidenced.
@@ 2 Data
The wind data sources in this article consist of official meteorological data from Met Éireann, citizen science or crowdsourced wind data in the form of personal weather stations, and reanalysis data from numerical weather simulation models. We exclusively focus on data from the Republic of Ireland, an island located in the east Atlantic. Its climate is strongly influenced by the ocean, with prevailing southwesterly to westerly winds, and wind strongest in coastal locations (Met Éireann, 2025b, ). An overview of each data source is provided in the following.
@@ 2.1 Met Éireann Data
As Ireland’s national meteorological service, Met Éireann maintains a network of quality-controlled weather stations, adhering to the standards set by the World Meteorological Organization (Oke,, 2018). Meteorological station data sourced from Met Éireann is regarded as ground truth in this study due to the high-quality measurement devices used and the optimal siting of instruments, ensuring no obstacles in the vicinity of measurement stations (Met Éireann, 2025a, ). Wind speed measurements are taken at a standard height of 10 m in open locations, ensuring no obstructions within a 100 m radius (Met Éireann, 2025b, ). The data is available in one minute intervals for 18 stations, and at hourly level for a further 5 stations. For consistency across all stations, including the PWS data introduced in Section 2.2, the data is aggregated to an hourly level by averaging sub-hourly measurements over the preceding hour. Aggregating high-frequency wind measurements is a standard practice in wind resource studies, as it balances temporal resolution with noise reduction and compatibility with other datasets (Torres et al.,, 2005). Figure 1 shows the location of the stations, and the median and inter-quartile range of wind speeds over a full year. Although there are only 23 weather stations which measure wind speeds in Ireland—a relatively small number—they are distributed across the country, ensuring good geographic coverage. In general, median wind speeds and wind speed variability are highest in coastal locations, particularly the west coast.
Wind speeds also exhibit a diurnal pattern, as illustrated in Figure 2, which presents the average hourly wind speeds over the year 2024 at four station locations. This shows higher wind speeds during the afternoon with lower wind speeds at night, and a stronger diurnal cycle at inland and leeward locations. This pattern is commonly seen in wind speeds and attributed to the effects of solar heating (Dai and Deser,, 1999).
@@
Figure 2: Hourly wind speed average over the year 2024 at four meteorological stations. The four sample stations are numbered on the map of median wind speeds in Figure 1.
@@ 2.2 Personal Weather Stations
The personal weather station (PWS) data is obtained from Weather Observation Website, 2011a , which provides measurements from stations set up by individual users. We selected this dataset because it is hosted on a single platform alongside meteorological station data and includes supporting information on station setup. Other sources of crowdsourced weather data include commercial platforms such as Weather Underground, (2025) and Netatmo, (2025), which could be considered in future research. The data from each station is sent digitally to the website which displays weather data and metadata on all stations. Unlike meteorological stations, there is no consistency of equipment or setup, nor is there any quality control. The website only provides information on the location in addition to a broad classification of the site attributes. These fall under three broad classifications, which are based on grading schemes from meteorological bodies (Weather Observation Website, 2011b, ), along with other stations which have not been classified. The three main classifications are described in Table 1:
@@
A Wind sensors calibrated within the last 10 years, mounted 10m above the ground on mast or pole, with no obstructions within 100m.
@@
Temporal resolution of the stations can vary, although they are generally at most hourly, and frequently contain data on a sub-hourly resolution, many of which are on a 5-minute resolution.
Figure 3 shows the distribution of wind speeds across different measurement sources, and highlights that wind speeds at PWS are generally lower when compared to meteorological stations, a similar finding to previous studies on crowdsourced data (e.g. Chen et al., 2021c, ). This may be due to the high dependency of wind measurements on the placement of equipment. Unlike weather variables such as temperature or air-pressure, local obstacles such as buildings, trees or hedges can have a significant effect on local wind speeds. Meteorological stations typically place anemometers at 10m heights, whereas the heights of the PWS measurements are assumed to be at less than 10m unless the station is in class A.
Figure 3: Violin plot showing the wind speed distribution for each group of stations in the Republic of Ireland. A, B, C, and U are the different classes of PWS, Met refers to meteorological weather stations. The black diamonds denote the median of each group. $n$ represents the number of stations in each class.
Consequently, directly including PWS wind speeds in a model, be it statistical or a physical weather model, would most likely lead to significant inaccuracies, in particular underestimation of wind speeds (Chen et al., 2021c, ). To highlight the potential usefulness of crowdsourced stations we select examples where we have a meteorological station and PWS in close proximity(<<15km) to each other, presented in Figure 4.
Figure 4: Location of meteorological stations in blue, and PWS in red. Four groups have been highlighted and sample time series for each group are shown in Figure 5.
@@
(a) 1. Mace Head (Met) and Glynsk (PWS, Class U). Pearson correlation = 0.86
(b) 2. Sherkin Island (Met) and Skibbereen (PWS, Class B). Pearson correlation = 0.76
(c) 3. Dublin Airport (Met) and Swords (PWS, Class C). Pearson correlation = 0.79
(d) 4. Oak Park (Met), Ballyhide (PWS, Class U). Pearson correlation = 0.83
@@
Figure 5: Comparison of wind speed time series recorded at nearby meteorological (Met) stations (in blue) and PWS (in red) at four locations around Ireland over the period July 13th - July 16th 2024. Panel captions indicate the Pearson correlation coefficient between the two time series.
@@
Although Figure 5 illustrates that PWS tend to underestimate the true wind speeds, there appears to be a strong correlation between the time series of the neighbouring stations. To compare measurements from nearby stations on a common scale, we compute the empirical cumulative distribution function (ECDF) for each station, and use this ECDF to transform wind speed time series into rank time series. Such time series express each observation as its rank (normalized in [0,1]) within that station’s ECDF. Figure 6 shows that after rank transformation neighbouring time series are in good agreement, and exhibit similar temporal behaviour and comparable ranks at each time point.
(a) 1. Mace Head (Met) and Glynsk (PWS, Class U). Pearson correlation = 0.85
(b) 2. Sherkin Island (Met) and Skibbereen (PWS, Class B). Pearson correlation = 0.78
(c) 3. Dublin Airport (Met) and Swords (PWS, Class C). Pearson correlation = 0.80
(d) 4. Oak Park (Met) and Ballyhide (PWS, Class U). Pearson correlation = 0.83
Figure 6: Comparison of wind speed rank time series recorded at nearby meteorological (Met) stations (in blue) and PWS (in red) at four locations around Ireland over the period July 13th - July 16th 2024. Note that the ECDFs used to compute the ranks cover the entire dataset while the plots display a three-day sample period. Panel captions indicate the Pearson correlation coefficient between the two rank time series, which is equivalent to the Spearman’s correlation coefficient between the raw wind speed time series.
@@
In general, PWS data appears to be correlated with nearby meteorological stations. However, as PWS can vary widely in reliability and data quality, we apply a series of quality assurance checks to the raw wind speed time series, and identify and remove stations that do not meet these checks. Bias correction methods for addressing remaining systematic differences are described in Section 3.
@@
First, we exclude PWS stations with excessive missing data. Only stations with at least 90% non-missing wind speed observations are retained, a criteria used in Lenzi and Genton, (2020). In addition, stations that do not upload data in near real-time are excluded because applications in renewable energy require timely updates.
@@
Second, we assess data quality using empirical Spearman’s rank correlations. Spearman’s correlation has been selected because it is robust to non-linear relationships, and therefore captures consistency in wind speed rankings rather than raw values, which is important when stations exhibit non-linear biases as illustrated in Figures 5 and 6. Figure 7 investigates how Spearman’s rank correlation is influenced by the distance between stations and the class of stations. As expected, meteorological stations display the highest pairwise correlations with each other, and the pairwise correlation decreases with the distance between the stations. Similarly, the trend of decreasing correlation with distance is seen for Met-PWS and PWS–PWS station pairs. Finally, one can notice that for a given separation distance the Met-Met stations are the more correlated, followed by Met-PWS and PWS-PWS.
@@
To remove unreliable stations, we extend the approach of Chen et al., 2021c , who discard stations with correlation below 0.5 relative to the nearest meteorological station. In our case, a PWS is retained only if at least five of its nearest neighbours (which may be meteorological or personal stations) exhibit a Spearman correlation greater than 0.5. This threshold is chosen as all meteorological stations meet this criteria, indicating it is a reasonable benchmark. While we do not correct or remove point-level anomalies in individual time series, this correlation based filtering removes stations that consistently report conflicting observations with nearby quality-controlled stations.